# Two dimensional dimers beyond planarity

### v2

@@ 6.2 Statement and proof of the Cauchy-Schwarz inequality
Theorem 6.3 (Cauchy-Schwarz inequality).
@@
[…]
@@
For the first two terms on the right side above we first note that $(\mu B_{x}-A_{x})_{x\in\mathbb{S}_{L,K}}$ is a reflection invariant vector of functions. Hence by Propositions 4.3 and 4.4 the assumptions in Lemma 6.2 are satisfied for the two point function $G(y)=\langle(\mu B_{o}-A_{o})(\mu B_{y}-A_{y})\rangle$, and we can then apply that lemma to conclude that
@@
$\geq\sum_{i=1}^{3}(G(5e_{i})-G(3e_{i}))\cdot\frac{1-\cos 3k_{i}}{2\sin^{2}k_{i}}\geq-c,$
by our assumptions on $k$. Note that for the last inequality we observed that, for any $x\neq o$, we have that $G(x)=\langle\sin(rs_{o}^{1})\sin(rs_{x}^{1})\rangle$, and used the positive goodness condition and Lemma 5.1 to upper bound $G(x)\leq$ $\langle\cos(rs_{o}^{1})\cos(rs_{x}^{1})\rangle$ $\leq c$ for any $x\in\mathbb{S}_{L,K}$ uniformly in $L$.
@@
Using again integration by parts and applying Proposition 5.4 we obtain that, for any $\beta>0$, if $\tilde{h}>0$ is not too large (depending on $\beta$), then there exists $c=c(\beta,\tilde{h})>0$ such that, for any $L,K\in 2\mathbb{N}\cup\{1\}$,
@@
where for the last inequality we also used the lower bound in Lemma 5.3 and the positive goodness for $\cos(rs_{o}^{1})$. Using again Proposition 5.4 for the three terms in (6.8) we then deduce that, under the same assumptions, $\frac{1}{4}\Delta\leq-\frac{c}{2}<0$ where $c$ is the same constant as in the previous display. Combining the estimates for the diagonal and the non-diagonal terms we then obtain that, under the assumptions of our theorem,
$\forall\mu\in\mathbb{R}\quad\langle(\mu B-A)\circ(\mu B-A)\rangle\geq-c$
@@ 6.3 Estimates for $\langle A\circ A\rangle$
We omit the subscripts in the expectation. First note the trivial bound $\sum_{k\in\mathbb{S}_{L,K}^{*}}\langle\sin^{2}(rs_{o}^{1})\rangle\leq\,|\mathbb{S}_{L,K}^{*}|\,\langle\sin^{2}(rs_{o}^{1})\rangle\leq cKL^{2}.$ Here we used the trigonometric identity $\sin^{2}x=\frac{1}{2}-\frac{1}{2}\cos(2x)$, and $\langle\cos(2rs_{o}^{1})\rangle\geq 0$ by the positive goodness condition. We now upper bound the off-diagonal term. For any $x\in\mathbb{S}_{L,K}$ we define the sets ${\mathcal{H}}_{x}=\mathcal{H}+x$, ${\mathcal{V}}_{x}=\mathcal{V}+x$, ${\mathcal{W}}_{x}=\mathcal{W}+x$, where the sum is with respect to the torus metric, we also recall the definitions $\mathbb{Z}^{M}_{odd}=(2\mathbb{Z}+1)\cap(-M/2,M/2]$, $\mathbb{Z}^{M}_{even}=2\mathbb{Z}\cap(-M/2,M/2]$, and define $\mathbb{Z}^{M}:=(-M/2,M/2]\cap\mathbb{Z}$ for any $M\in\mathbb{N}$. Using translation invariance we have that
@@
$+\sum_{x\in\mathbb{S}_{L,K}}\sum_{y\in\mathcal{W}_{x}}\cos(k\cdot(x-y))\langle\sin(rs_{x}^{1})\sin(rs_{y}^{1})\rangle$
$\leq 2KL\sum_{x=(-\frac{L}{2}+1)e_{1}}^{\frac{L}{2}e_{1}}\sum_{y\in\mathbb{Z}^{L}_{odd}e_{1}+x}\cos(k\cdot(x-y))\langle\sin(rs_{x}^{1})\sin(rs_{y}^{1})\rangle$
@@
$+\sum_{y\in\mathcal{W}}\cos(k\cdot y)\langle\sin(rs_{o}^{1})\sin(rs_{y}^{1})\rangle\biggr{)}$
$\quad\leq C^{\prime}KL^{2},$
@@
$\sum_{k\in\mathbb{S}_{L,K}^{*}\setminus\mathbb{S}_{L,K,+}^{*}}\langle A\circ A\rangle_{L,K,\beta,h,_{L,K},r}\geq-c\,K\,L^{2}$
@@
$-cKL\Biggl{(}\sum_{y\in\mathbb{Z}^{L}_{odd}e_{1}}\langle\sin(rs_{o}^{1})\sin(rs_{y}^{1})\rangle+\sum_{y\in\mathbb{Z}_{odd}e_{2}}\langle\sin(rs_{o}^{1})\sin(rs_{y}^{1})\rangle\Biggr{)}\\ \geq-cKL^{2}(\langle\sin(rs_{o}^{1})\sin(rs_{e_{1}}^{1})\rangle+\langle\sin(rs_{o}^{1})\sin(rs_{e_{2}}^{1})\rangle)\geq-c^{\prime}KL^{2},$ (6.14)
@@
7 Mermin-Wagner Theorem
We deduce our Mermin Wagner theorem from the following upper bound for the magnetisation.
Proposition 7.1 (Magnetisation bound).
Suppose that $\beta\geq 0$, $r\in\{1,2\}$, and that $\tilde{h}>0$ is small enough. There a constant $c=c(\beta)<\infty$ such that, for any $L,K\in 2\mathbb{N}\cup\{1\}$ satisfying $K\leq(\log L)^{\frac{1}{2}}$, we have
$m_{\mathbb{S}_{L,K}}(\beta,h_{L,K},r)\leq c\sqrt{\frac{K}{\log L}},$
@@ 7 Mermin-Wagner Theorem
Let us now define the set $\mathcal{U}:=\{(L,K)\in(2\mathbb{N}+1)^{2}\,:\,K\leq\sqrt{\log L}\}$. Suppose that for any $\varphi>0$ we have that
$m_{\mathbb{S}_{L,K}}(\beta,h_{L,K},r)\geq\varphi\sqrt{\frac{K}{\log L}}\,\,\mbox{ for infinitely many choices of $(L,K)\in\mathcal{U}$.}$ (7.2)
We will find a contradiction with this assumption, deducing that there exists a value $\varphi>0$ such that $m\geq\varphi\sqrt{\frac{K}{\log L}}$ can only hold for finitely many choices of $(L,K)\in\mathcal{U}$. This implies the claim of the proposition, as desired.
@@
$\frac{1}{|\mathbb{S}_{L,K}|}\sum_{k\in\mathbb{S}_{L,K,+}^{*}}\frac{\varphi^{2}}{c_{2}|k|^{2}+\frac{c_{1}}{|\mathbb{S}_{L,K}|}}\leq c_{3}\frac{\log L}{K},$ (7.3)
@@
$c_{4}\varphi^{2}\frac{\log(L)}{K}\leq c_{3}\frac{\log L}{K},$ (7.4)
@@
We now state our main theorem in the probabilistic framework of random loops with arbitrary weight function.
Theorem 7.2 (Mermin-Wagner theorem for double dimers).
@@
$\sum_{x\in\mathbb{S}_{L,K}^{o}}\frac{\mathcal{G}_{L,K,\beta}^{(1)}(o,x)}{|\mathbb{S}^{o}_{L,K}|}$ $\leq c\sqrt{\frac{K}{\log L}}$ (7.5)
$\sum_{x\in\mathbb{S}_{L,K}}\frac{\mathcal{G}_{L,K,\beta}^{(\ell)}(o,x)}{|\mathbb{S}_{L,K}|}$ $\leq c\sqrt{\frac{K}{\log L}},$ (7.6)
where the second inequality holds only if $U$ is nice when $\ell=1$ and only if $U$ is fast decaying when $\ell=2$.
@@
Proof of Theorem 1.1 when $N=1$ and of Theorems 1.2, 2.1, and 7.2.
Fix $\beta>0$, $\tilde{h}>0$ small enough. We apply the first claim in Proposition 5.4 and Proposition 7.1 to deduce our main theorem for the spin system, Theorem 2.1.
@@
Since for the monomer double dimer model $U$ is fast decaying (and nice), when $\beta>0$ is finite we can apply Theorem 7.2 with $\ell=2$ and (3.6) to deduce our main theorem on the expected length of the loops in the monomer double dimer model, Theorem 1.1, when $N=1$ and $\rho=1/\beta>0$.
Moreover, we use (7.5) (which does not require the assumption that $U$ is nice) for $U=\mathbbm{1}_{\{n=1\}}$, corresponding to the double dimer model, (3.7), and the fact that $\mathcal{C}(o,x)=0$ if $x\in\mathbb{S}_{L,K}^{e}$ to obtain our second main theorem on monomer correlations in the dimer model, Theorem 1.2.
In order to obtain Theorem 1.1 when $N=1$ and $\rho=0$ we use that
$\mathbb{P}_{L,K,1,0}(x\leftrightarrow y)=\sum\limits_{z\sim x,k\sim y}\,\mathcal{C}(x,y)\mathcal{C}(z,k)\leq d^{2}\,\mathcal{C}(x,y),$
@@
where the identity follows from the fact that by [21, Figure 2] $\mathcal{C}(x,y)\mathcal{C}(z,k)$ corresponds to the probability that in the double dimer model a loop connects $x$ to $y$ with the blue dimers of such loop touching $x$ and $y$ being on the edges $\{x,z\}$ and $\{y,k\}$, while the inequality follows from the fact that $\mathcal{C}(z,k)\leq 1$ for any $z,k\in\mathbb{S}_{L,K}$ by [23, equation (2.25)]. Hence using the previous inequality and Theorem 1.2 we deduce Theorem 1.1 when $\rho=0$ and $N=1$.
7.1 Multiple dimer model
In order to treat the case of integer $N>1$ we need to generalise the spin representation considering $S=[0,2\pi)^{2N}$ as a local spin space. For each $\boldsymbol{s}\in S^{V}$, the Hamiltonian can be generalised to,
$H(\boldsymbol{s})=\sum\limits_{u=1}^{2N}\sum\limits_{\{i,j\}\in E}e^{i(s_{i}^{u}-s_{j}^{u})}+\,h\sum\limits_{k\in V}\,2\,\cos(rs^{1}_{k}),$
and the function $g$ has to be generalised as follows,
$g(s)=\frac{1}{(2\pi)^{2N}}\sum\limits_{n=(n_{1},\ldots n_{N})\in\mathbb{N}_{0}^{N}}U(n_{1},\dots,n_{N})e^{-\sum_{u=1}^{N}n_{u}\,i\,(s^{2u-1}-s^{2u})}$
@@ 7.1 Multiple dimer model
where $U:\mathbb{N}^{N}\mapsto\mathbb{R}$ is the weight function and $s=(s^{1},\ldots,s^{N})$.
The double dimer representation which is obtained by applying the strategy of Proposition 8 below leads to a double dimer representation which is analogous to the one which has been introduced in Section 3, the only difference is that dimers have $2N$ possible colours, $1,\ldots,2N$, and at each vertex the number of dimers of odd colour $i\in[1,2N]$ equals the number of dimers of colour $i+1$. In this representation, each vertex touched by $n_{i}$ dimers of colour $2i$ (which is then also touched by $n_{i}$ dimers of colour $12i-$ 1) for each $i\in\{1,\ldots,N\}$ receives a multiplicative weight $U(n_{1},\ldots n_{N})$.
Pairings can be introduced analogously to Section 3 to obtain open and closed path, namely at each vertex $x\in V$ dimers of odd colour $i$ can only be paired to paths of colour $i+1$ and viceversa. This leads to a random walk loop soup in which each close path is a sequence of dimers whose colours are $2i$, $12i-$ 1, $2i$, $12i-$ 1, $\ldots$,… , or $12i-$ 1, $2i$, $12i-$ 1, $2i$, $\ldots$ for some integer $i\in[0,N]$.
We obtain the monomer double dimer-model (1.1) with arbitrary $N\in\mathbb{N}$ in the special case,
$U(n_{1},n_{2},\ldots n_{N})=\begin{cases}1&\mbox{ if $n_{1}+\ldots+n_{N}\in\{0,1\}$,}\\ 0&\mbox{ otherwise.}\end{cases}$
Indeed, in this case the local time of each vertex is either zero or one, the same as in (1.1), and each loop of length two has multiplicity $N$ (we have $N$ possibilities for choosing whether the dimers constituting this loop have colours $1$ and $2$, $3$ and $4$, $\ldots$, or $12N-$ 1 and $N$) while each loop of length greater than two has multiplicity $2N$ (with each case corresponding to one of the possible sequences $2i$, $12i-$ 1, …$2i$ or $12i-$ 1, $2i$, …$12i-$ 1). This is in correspondence to (1.1), in which each loop of length two has multiplicity $1$ and each loop of length greater than two has multiplicity $2$ and, in addition, there is a multiplicative weight $N$ for each loop.
8 Proof of Proposition 3.1
Proof.
We observe that the factor $e^{\beta H(\boldsymbol{s})}$ in (2.4) can be seen as a product of exponential functions. Each such exponential function is associated to an edge of the extended graph and to a colour. Recall the definition of $G=(V,E)$ , of $G_{en}=(V_{en},E_{en})$, $E_{s}$, and $E_{g}$. Assume that $A\subset V$. We start by performing a Taylor expansion for each such exponential in the product, obtaining
$\int_{S^{V}}\boldsymbol{ds}\,\boldsymbol{g}(\boldsymbol{s})\prod_{z\in A}g_{u,z}(s_{z})\,\,e^{\beta H(\boldsymbol{s})}\,\prod_{x\in V^{e}}e^{i(u^{1}_{x}s_{x}^{1}-u^{2}_{x}s_{x}^{2})}\prod_{x\in V^{o}}e^{-i(u^{1}_{x}s_{x}^{1}-u^{2}_{x}s_{x}^{2})}\\ =\sum\limits_{m^{1}\in\mathbb{N}_{0}^{E}}\prod_{e\in E}\frac{\beta^{m^{1}_{e}}}{m^{1}_{e}!}\sum\limits_{m^{2}\in\mathbb{N}_{0}^{E}}\prod_{e\in E}\frac{\beta^{m^{2}_{e}}}{m^{2}_{e}!}\sum\limits_{m^{g}\in\mathbb{N}_{0}^{E_{g}}}\prod_{e\in E_{g}}\frac{(\beta h)^{m^{g}_{e}}}{m^{g}_{e}!}\\ \times\int_{S^{V}}\boldsymbol{ds}\,\boldsymbol{g}(\boldsymbol{s})\prod_{z\in A}g_{u,z}(s_{z})\,\,\prod_{x\in V^{e}}e^{i(u^{1}_{x}s_{x}^{1}-u^{2}_{x}s_{x}^{2}+\,\tilde{n}_{x}(m^{1})s_{x}^{1}\,-\tilde{n}_{x}(m^{2})s_{x}^{2})}\prod_{x\in V^{o}}e^{-i(u^{1}_{x}s_{x}^{1}-u^{2}_{x}s_{x}^{2}+\,\tilde{n}_{x}(m^{1})s_{x}^{1}\,-\tilde{n}_{x}(m^{2})s_{x}^{2})}\\ \times\prod_{x\in V}(2\cos(rs_{x}^{1}))^{m_{\{x,g\}}^{g}},$ (8.1)
where $\tilde{n}_{x}(m^{i})=\sum_{y\sim x}m_{\{x,y\}}^{i}$, recalling that $\sim$’ is the neighbourhood relation in $G$ and not in $G_{en}$. We interpret $m_{\{x,y\}}^{i}$ as the number of dimers of colour $i$ on $\{x,y\}$ and $m_{\{x,g\}}^{g}$ as the number of dimers on $\{x,g\}$ (independently from their colour). Given realisation of the vector $m^{g}$ in the previous sum, we perform the following expansions at each vertex

### v3

@@ 6.2 Statement and proof of the Bogoliubov-type inequality
Theorem 6.3 (Bogoliubov-type inequality).
@@
Here and in the next proofs we omit the subscripts from $\langle\,\,\,\rangle$ to simplify the notation. Take $\mu\in\mathbb{R}$, using bilinearity of the inner product,
@@
For the first two terms on the right side above we first note that $(\mu B_{x}-A_{x})_{x\in\mathbb{S}_{L,K}}$ is a reflection invariant vector of functions. Hence by Propositions 4.7 and 4.8 the assumptions in Lemma 6.2 are satisfied for the two point function $G(y)=\langle(\mu B_{o}-A_{o})(\mu B_{y}-A_{y})\rangle$, and we can then apply that lemma to conclude that
@@
$\geq\sum_{i=1}^{3}(G(5e_{i})-G(3e_{i}))\cdot\frac{1-\cos(3k_{i})}{2\sin^{2}(k_{i})}\geq-c,$
by our assumptions on $k$. Note that for the last inequality we observed that, for any $x\neq o$, we have that $G(x)=\langle\sin(rs_{o}^{1})\sin(rs_{x}^{1})\rangle$, and used the positive-expectation condition and Lemma 5.1 to upper bound $G(x)\leq$ $\langle\cos(rs_{o}^{1})\cos(rs_{x}^{1})\rangle$ $\leq c$ for any $x\in\mathbb{S}_{L,K}$ uniformly in $L$.
@@
Using again integration by parts and applying Proposition 5.5 we obtain that, for any $\beta>0$, if $\tilde{h}>0$ is small enough (depending on $\beta$), then there exists $c=c(\beta,\tilde{h})>0$ such that, for any $L,K\in 2\mathbb{N}\cup\{1\}$,
@@
where for the last inequality we also used the lower bound in Lemma 5.3 and the positive-expectation for $\cos(rs_{o}^{1})$. Applying again (5.11) for the three terms in (6.8) (which implies, in particular, $m\leq c\tilde{h}$ and $\langle\cos(2rs_{o}^{1})\rangle)\leq c\tilde{h}$), we then deduce that, under the same assumptions, $\frac{1}{4}\Delta\leq-\frac{c}{2}<0$ where $c$ is the same constant as in the previous display. Therefore, the diagonal term $\langle(\mu B_{o}-A_{o})^{2}\rangle\geq 0$.
Combining the estimates for the diagonal and the non-diagonal terms we then obtain that, under the assumptions of our theorem,
$\forall\mu\in\mathbb{R}\qquad\langle(\mu B-A)\circ(\mu B-A)\rangle\geq-c$
@@ 6.3 Estimates for $\langle A\circ A\rangle$
We omit the subscripts in the expectation. First note the trivial bound
$\sum_{k\in\mathbb{S}_{L,K}^{*}}\langle\sin^{2}(rs_{o}^{1})\rangle\leq\,|\mathbb{S}_{L,K}^{*}|\,\langle\sin^{2}(rs_{o}^{1})\rangle\leq cKL^{2}.$
Here we used the trigonometric identity $\sin^{2}x=\frac{1}{2}-\frac{1}{2}\cos(2x)$, and $\langle\cos(2rs_{o}^{1})\rangle\geq 0$ by the positive-expectation condition. We now upper bound the off-diagonal term. For any $x\in\mathbb{S}_{L,K}$ we define the sets ${\mathcal{H}}_{x}=\mathcal{H}+x$, ${\mathcal{V}}_{x}=\mathcal{V}+x$, ${\mathcal{W}}_{x}=\mathcal{W}+x$, where the sum is with respect to the torus metric, we also recall the definitions $\mathbb{Z}^{M}_{odd}=(2\mathbb{Z}+1)\cap(-M/2,M/2]$, $\mathbb{Z}^{M}_{even}=2\mathbb{Z}\cap(-M/2,M/2]$, and define $\mathbb{Z}^{M}:=(-M/2,M/2]\cap\mathbb{Z}$ for any $M\in\mathbb{N}$. Using translation invariance we have that
@@
$+\sum_{x\in\mathbb{S}_{L,K}}\sum_{y\in\mathcal{W}_{x}}\cos(k\cdot(x-y))\langle\sin(rs_{x}^{1})\sin(rs_{y}^{1})\rangle.$
Now we see that the previous quantity is less or equal to
$2KL\sum_{x=(-\frac{L}{2}+1)e_{1}}^{\frac{L}{2}e_{1}}\sum_{y\in\mathbb{Z}^{L}_{odd}e_{1}+x}\cos(k\cdot(x-y))\langle\sin(rs_{x}^{1})\sin(rs_{y}^{1})\rangle$
@@
$+\sum_{y\in\mathcal{W}}\cos(k\cdot y)\langle\sin(rs_{o}^{1})\sin(rs_{y}^{1})\rangle\biggr)\leq C^{\prime}KL^{2},$
@@
$\sum_{k\in\mathbb{S}_{L,K}^{*}\setminus\mathbb{S}_{L,K,+}^{*}}\langle A\circ A\rangle_{L,K,\beta,h_{L,K},r}\geq-c\,K\,L^{2}$
@@
$cKL\Biggl(\sum_{y\in\mathbb{Z}^{L}_{odd}e_{1}}\langle\sin(rs_{o}^{1})\sin(rs_{y}^{1})\rangle+\sum_{y\in\mathbb{Z}^{L}_{odd}e_{2}}\langle\sin(rs_{o}^{1})\sin(rs_{y}^{1})\rangle\Biggr)\\ \geq-cKL^{2}(\langle\sin(rs_{o}^{1})\sin(rs_{e_{1}}^{1})\rangle+\langle\sin(rs_{o}^{1})\sin(rs_{e_{2}}^{1})\rangle)\geq-c^{\prime}KL^{2},$ (6.14)
@@
7 Mermin–Wagner Theorem
The Mermin–Wagner theorem states that in two dimensions the magnetisation vanishes when the external field is taken to zero, independently of the inverse temperature. This holds for spin systems with continuous symmetries, but not, for instance, for the Ising model in $d>1$ at low temperature. We now state our analogue of the Mermin–Wagner theorem for our general framework. In contrast to the classical setting, where one first takes the thermodynamic limit and then sends the external field to zero, here, for technical reasons, we must let the field go to zero and the volume diverge simultaneously, with a suitable scaling.
Theorem 7.1 (Mermin–Wagner theorem).
Suppose that $\beta\geq 0$, $r\in\{1,2\}$, and that $\tilde{h}>0$ is small enough. There a constant $c=c(\beta)<\infty$ such that, for any $L,K\in 2\mathbb{N}\cup\{1\}$ satisfying $K\leq(\log(L))^{\frac{1}{2}}$, we have
$m_{\mathbb{S}_{L,K}}(\beta,h_{L,K},r)\leq c\sqrt{\frac{K}{\log(L)}},$
@@ 7 Mermin–Wagner Theorem
Let us now define the set $\mathcal{U}:=\{(L,K)\in(2\mathbb{N}+1)^{2}\,:\,K\leq\sqrt{\log(L)}\}$. Suppose that for any $\varphi>0$ we have that
$m_{\mathbb{S}_{L,K}}(\beta,h_{L,K},r)\geq\varphi\sqrt{\frac{K}{\log(L)}}\,\,\mbox{ for infinitely many $(L,K)\in\mathcal{U}$.}$ (7.2)
We will find a contradiction with this assumption, deducing that there exists a value $\varphi>0$ such that $m\geq\varphi\sqrt{\frac{K}{\log(L)}}$ can only hold for finitely many choices of $(L,K)\in\mathcal{U}$. This implies the claim of the proposition, as desired.
@@
$\frac{1}{|\mathbb{S}_{L,K}|}\sum_{k\in\mathbb{S}_{L,K,+}^{*}}\frac{\varphi^{2}}{c_{2}|k|^{2}+\frac{c_{1}}{|\mathbb{S}_{L,K}|}}\leq c_{3}\frac{\log(L)}{K},$ (7.3)
@@
$c_{4}\varphi^{2}\frac{\log(L)}{K}\leq c_{3}\frac{\log(L)}{K},$ (7.4)
@@
The next theorem shows how the decay of magnetisation implies the decay of the two–point functions.
Theorem 7.2 (From magnetisation decay to correlations decay).
@@
$\sum_{x\in\mathbb{S}_{L,K}^{o}}\frac{\mathcal{G}_{L,K,\beta}^{(1)}(o,x)}{|\mathbb{S}^{o}_{L,K}|}$ $\leq c\sqrt{\frac{K}{\log(L)}}$ (7.5)
$\sum_{x\in\mathbb{S}_{L,K}}\frac{\mathcal{G}_{L,K,\beta}^{(\ell)}(o,x)}{|\mathbb{S}_{L,K}|}$ $\leq c\sqrt{\frac{K}{\log(L)}},$ (7.6)
where the second inequality holds only if $U$ is normalized monotone when $\ell=1$ and only if $U$ is fast decaying when $\ell=2$.
@@
Proof of Theorems 1.1, 2.2, and 7.2.
Fix $\beta>0$, $\tilde{h}>0$ small enough. We apply the first claim in Proposition 5.5 and Theorem 7.1 to deduce our main theorem for the spin system, Theorem 2.2.
@@
We now use our Theorem 7.2 to deduce our main theorems on the dimer and monomer double-dimer model, respectively Theorems 1.1 and 1.2.
For Theorem 1.1, recall that if we choose $U=\mathbbm{1}_{\{n=1\}}$ then the two-point function $\mathcal{G}^{(1)}_{L,K,\beta}(o,x)$ corresponds to the monomer-monomer correlation of the dimer model, $\mathcal{C}(o,x)$, as we observed in (3.10). Hence, the first inequality in Theorem 7.2, together with the observation that $\mathcal{C}(o,x)=0$ if $x\in\mathbb{S}_{L,K}^{e}$, implies directly Theorem 1.1, as desired.
We now prove Theorem 1.2. Note that the second claim, (1.6), follows from the first claim, (1.5), by a direct application of the Markov inequality. We choose $U(0)=U(1)=1$ and $U(n)=0$ for all $n>1$, giving the monomer-double dimer model with monomer activity $\rho=1/\beta$ when $\beta>0$, as we observed in Section 3.2. Since under this choice $U$ is fast decaying (and normalized monotone), we can apply Theorem 7.2 with $\ell=2$ and $\ell=1$ and (3.7) to deduce (1.5) and (1.7) in Theorem 1.2 when $\rho=1/\beta>0$.
When $\rho=0$, (1.7) in Theorem 1.2 follows directly from Theorem 1.1 since in this case $\mathbb{G}_{L,K,0}(o,x)=\mathcal{C}(o,x)$, as we observed in Section 3.2.
In order to obtain (1.5) in Theorem 1.2 when $\rho=0$ we use that
$\mathbb{P}_{L,K,0}(x\leftrightarrow y)=\sum\limits_{z\sim x,k\sim y}\,\mathcal{C}(x,y)\mathcal{C}(z,k)\leq d^{2}\,\mathcal{C}(x,y),$
@@
where the identity follows from the fact that, by [29, Figure 2], $\mathcal{C}(x,y)\mathcal{C}(z,k)$ corresponds to the probability that in the double-dimer model a loop connects $x$ to $y$ with the blue dimers of such loop touching $x$ and $y$ being on the edges $\{x,z\}$ and $\{y,k\}$, while the inequality follows from the fact that $\mathcal{C}(z,k)\leq 1$ for any $z,k\in\mathbb{S}_{L,K}$ by [33, equation (2.25)]. Hence using the previous inequality and Theorem 1.1 we deduce that
$\sum\limits_{x\in\mathbb{S}_{L,K}}\frac{\mathbb{P}_{L,K,0}(o\leftrightarrow x)}{|\mathbb{S}_{L,K}|}\leq cd^{2}\sqrt{\frac{K}{\log(L)}},$ (7.7)
from which we deduce (1.5).
8 Appendix
8.1 Multiple-type dimer model
We now explain how to extend our framework to the case of dimers of $2N$ different types, where $N$ is an integer. The local spin space is the set $\Xi=[0,2\pi)^{2N}.$ The configuration space is the set $\Omega_{s}=\Xi^{V}$, with $\boldsymbol{s}=(s_{z})_{z\in V}\in\Omega_{s}$, and $s_{z}=(s_{z}^{1},\ldots,s_{z}^{2N})$. The spin variable at $z$ is then a vector
$S_{z}=(S_{z}^{1},\ldots,S_{z}^{2N}),$
@@ 8.1 Multiple-type dimer model
where the function $S_{z}^{k}:\Omega_{s}\mapsto\mathbb{C}$ is defined for each $k\in\{1,\ldots,2N\}$ as